0.5 _____ 0.7 = 0.35
 0.5 _____ 0.7 = 0.35
 0.5 _____ 0.7 = 0.35

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

0.5 × 0.7 = 0.35
While multiplying two decimal numbers, first multiply them as whole numbers.
Countthe number of digits to the right of the decimal point in both the decimal numbers. Add the number of digits counted. Put the decimal point in the product by counting the number of digits equal to sum obtained from its rightmost place.
= 0.5 × 0.7 =5 × 7
= 35
= 0.35


Thanks